The supply curve for a certain industry is given by S(p) = p−3 and demand by D(p) = 33 − 3p.

5. A tax is levied at the rate of $2 per unit. What are the new buyer and seller prices, pb, ps, and the new quantity q? Which price changed more?

6. What is the total tax revenue and the producer and consumer surplus under this tax? What is the deadweight loss (DWL)?

7. If the tax doubles from $2 to $4 per unit, by what factor does the DWL change? (It is possible to answer this question without recalculating the DWL, by geometric intuition regarding the DWL triangle.)

8. Suppose, before any tax was passed, the supplier believed the market would not change any time soon, and locked in labor contracts which commit irrevocably to the current (no-tax) quantity of supply. What would the new supply curve be? Draw it and write its equation. 1 2

9. With the new supply curve as in (8), who would bear a $2 per unit tax? What would the deadweight loss be?
Answer #1

s(P P-3 DCP)- B3-2p T: 42/uaih NTo Tar: equitibrun, sCP) DC A+ P-3-33- Bp 36 -4p . 9=p- 3 9-3 -G 2 No PPsT DlP)- 33-3 2-3 -3Total Sumplus rithont tan - bx-2-6 PS= X ab: 18 Total lunplus 9t691 24 CSt PS with Tan: Cst Ps+ gos. Revenu 3. 37S t 10 125 tfixat Auppuer mupply wiu beuou at curent lenel then vahcal and t crve would not yafteet ehange uuth phanges in price Cafe tam